

********************************************************************
UARS WEEKLY STATUS REPORT - 20 JUL 2001
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, taking solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. We continue to gather enough data to meet all the SOLSTICE
science objectives during this period of reduced tape recorder operations.
***********
SUSIM
***********
SUSIM had a commanding error which caused it to miss two MgII index
scans. This now has been corrected. Otherwise, SUSIM has completed
its daily solar scans this past week during periods of predicted UARS
telemetry.
***********
ACRIM
***********
Following a loss of power in early May, the ACRIM instrument has not
responded to power-up commands. The ACRIM team is working on a scenario of
power-ups and downs to test the apparent inability of the instrument power
supply to activate.
The ACRIM2 total solar irradiance database has been re-processed and is
complete up to the May 2001 power loss. The new processing scheme utilizes
improved algorithms developed for the ACRIMSAT/ACRIM3 experiment whose
science mission began in April 2000. Reprocessing has found additional
results and more degradation comparisons that have resulted in a much
improved version of the total solar irradiance results from the experiment.
These results are available at the ACRIM Science Team web site
(http://acrim.com) and at NASA's LaRC and GSFC DAAC's.
***********
PEM
***********
***** ALL PEM COMMUNICATION SHOULD NOW BE THROUGH THE PEM2K:: NODE *****
The PEM team reports that the PEM:: node computer has failed and has
been taken off line. PEM:: appears to have hardware problems that take the
system down every few hours. Automatic transfer NRT tests to PEM2K from
both CDHF2 and CDHFB have succeed and the NRT files have been verified against
those captured by PEM, which confirm the NRT process is correct. Thus, PEM
confirms NRT data from CDHFB. Thanks to Steve Bove for his work assisting
PEM with this process.
PEM/AXIS was activated in activated in full power mode on July 3. At
the current time, AXIS is not generating energy deposition. SwRI is looking
into this. Tests of AXIS production code on the rehost computer are continuing.
At the present time, PEM energy deposition from its spectrometers is not
being generated. Recent data configuration changes have affected PEM software
causing multiple failures. These failures are being reviewed by the PEM team.
In order to correct for new telemetry requirements, PEM will need to update
processing code with a patch delivery. The delivery patch may include changes
which would allow the PEM software to execute on a DEC Alpha machine. GSFC is
helping PEM with some DEC Alpha particulars which have caused PEM level 3
code to fail on DEC Alpha machines (software below level 3 executes on the
DEC Alpha).
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
On 17 July 2001 the MPG participated in the biweeekly
three-TDRS meeting. These biweekly meetings are
being held to monitor the development of software
needed to allow UARS to begin using three TDRS
satellites for commanding and data acquistion.
Daily processing of instrument Daily Activity Plans
continues with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 07 July 2001 (Orbit 53686, GMT Day 188) through 13 July 2001
(Orbit 53790, GMT Day 194).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 194/0047 53776 Power OFF
WINDII 194/0643 53780 Power ON
The beta angle ranged between the angles of 33.7 degrees to 59.0 degrees for
this report period. The beta angle is now increasing towards a maximum angle
of 78 degrees on 20 July 2001 (DOY 201).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.2 to 1.6 degrees C. Additional battery performance data for each
battery on 6 July 2001 and 13 July 2001 are provided in Appendix D.
The clock error ranged from Ð11.8 to 10.5 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
¥ OUT-OF-LIMITS = 00
¥ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
¥ FOT = 19
¥ HALOE = 01
¥ WINDII = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
¥ FOT = 02
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/316 - 2287.496493
· Transponder B: 01/119 - 2287.499116
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 01/193 show the observatory data loss to be 302 hours, 44
minutes, 57.624 seconds. This shows a zero loss since the last report period.
This is a 0.39 percent data loss, which equals a 99.61 percent data capture for
the mission.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
An instrument anomaly was experienced on 05 May 2001. The instrument involved
was the ACRIM II instrument. All ACRIM telemetry for currents and voltages all
dropped to zero values suddenly. The ACRIM team in coordination with the
UARSFOT conducted analysis and command tests to determine the cause of the
anomaly. Analysis shows that the anomaly was caused by failed circuitry in the
ACRIM electronics box. The FOT is awaiting a future plan of action from the
ACRIM team.
The software test team, in conjunction with the FOT, has completed a large
portion of the testing for the new front end processor. The new TAC still has
yet to undergo DSN and realtime TDRS testing, but has completed simulations of
receipt of telemetry, commanding, and NCC/GCMR testing. DSN testing is
scheduled for 18 July 2001.
UARSFOT will also begin supporting testing of the rehosted UARS ground system.
A prototype system has been built in accordance with the UARS documentation and
operational database. The demo of the prototype system will be held on 02
August 2001
UARS attitude has been adjusted over the course of the last few yaw cycles by
loading new alignment matrices for the TAM. Good alignments have been found
for both directions of flight and the performance of the TAM should remain
stable and accurate.
MEETINGS
The FOT held the UARS Three TDRS Kickoff Meeting on 03 July 2001. This meeting
was established to track the progress and resolve problems of all entities
involved in the implementation of the three TDRS mode. This meeting will
become a biweekly meeting until the three TDRS mode is operational.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 08 August 2001 (DOY 220). This
maneuver will be a Forward to Reverse Flight maneuver (Southern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 13 July 2001
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
53690 188/0740 01-445 NBTR NOOPS D-188
53697 188/1759 01-446 NBTR NOOPS D-189
53704 189/0600 01-446 NBTR NOOPS D-189
53711 189/1650 01-446 NBTR NOOPS D-189
53712 189/1800 01-447 NBTR NOOPS D-190
53719 190/0600 01-447 NBTR NOOPS D-190
53727 190/1801 01-447 NBTR NOOPS D-190
53727 190/1801 01-448 NBTR NOOPS D-191
53729 190/2120 01-449 OBC/CLK Clk Adj 0.0 to 20.6 msec/day
53740 191/1506 01-450 OBC TAM Alignment
53742 191/1801 01-448 NBTR NOOPS D-191
53750 192/0741 01-451 NBTR NOOPS D-192
53753 192/1143 01-451 NBTR PLAYBACK/ /RTS-58
53753 192/1143 01-452 OBC/CLK Clk Adj 20.6 to 0.0 msec/day
53756 192/1620 01-454 TAM TAM ALIGNMENT
53757 192/1802 01-453 NBTR NOOPS D-193
53765 193/0642 01-453 NBTR NOOPS D-453
53776 194/0047 01-455 HALOE Power OFF
53780 194/0643 01-456 WINDII Power ON
53782 194/1005 01-457 OBC Clk Adj 0.0 to 20.6 msec/day
53785 194/1438 01-458 OBC TAM Alignment
______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 13 July 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 13 July 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
53729 190/2128 Clk Adj 0.0 to 20.6 msec/day
53753 192/1150 Clk Adj 20.6 to 0.0 msec/day
53782 194/1008 Clk Adj 0.0 to 20.6 msec/day
_______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 13 July 2001
6 July 2001 - GMT Day 187 Beta = 29.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- -------- ---------- ------ ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/+0.0 12.0 0.000 0.0
2 +11.2/-134.4 +5.79/+4.51 +28.4/-14.4 24.8 1.030 20.7
3 +22.4/-56.0 +4.83/+3.23 +28.0/-15.2 24.8 1.020 20.7
13 July 2001 - GMT Day 194 Beta = 59.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- -------- ---------- ------ ------- ------
1 +728.0/+728.0 -3.68/-4.94 +0.0/+0.0 12.0 0.000 0.0
2 +11.2/-33.6 +3.87/+2.60 +22.4/-14.4 26.4 1.030 18.5
3 +22.4/-22.4 +2.92/+1.66 +21.6/-14.4 26.4 1.020 18.0
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ---------
-no changes V/T 5
_______________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 13 July 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-042 FOT 53697 188/1759Z N/A Data dropouts/NISN router switch
01-043 FOT 53722 190/0920Z 23951 6 Second TLM dropout at WS
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-042 FOT 53697 188/1759Z N/A Data dropouts/NISN router switch
01-043 FOT 53722 190/0920Z 23951 6 Second TLM dropout at WS
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-041 FOT see list July 01 14170 Generic Late Acquisition
SEVNTFW01/28